# Open Data from LIGO, Virgo, and KAGRA through the First Part of the Fourth Observing Run

### v2

@@ 2.1 Observing Time
The instruments operate in a series of observing runs described in GWTC:Introduction, with the following dates:
@@ 4 Data Quality
Noise in the strain data is sometimes approximated as stationary and Gaussian. However, real strain data are also affected by disturbances originating from instrumental and environmental sources (LIGOScientific:2019hgc; LIGO:2024kkz). These artifacts can manifest as broadband, non-Gaussian, short-duration features known as glitches (Nuttall:2018xhi; Glanzer:2022avx), or as narrowband spectral features referred to as spectral lines (Covas_2018). In addition, GW observatories can experience periods of degraded performance due to issues in control systems, calibration procedures, or environmental disturbances. These episodes can compromise data quality over extended durations, reducing the sensitivity of the detectors and potentially biasing astrophysical parameter estimation (Pankow:2018qpo; Macas:2022afm; Kwok:2021zny; Ghonge:2023ksb). To support robust analyses, LVK employs a suite of data-quality metrics designed to identify and allow mitigation of the impact of instrumental and environmental noise. An overview of LIGO detector characterization efforts is available in LIGO:2024kkz. Detector characterization methods for past observing runs are described in LIGO:2021ppb; Virgo:2022ysc; Nuttall:2015dqa; LIGOScientific:2016gtq; LIGOScientific:2017tza; 2018PhRvD..97h2002C; 2021CQGra..38n5001N; Virgo:2022kwz; and KAGRA:detchar.
@@ 4.1 Hardware Injections
Hardware injection times are available on the GWOSC website as Timelines, with additional notes in the documentation for each data release. In O4a there were no injections labeled CBC, BURST, DETCHAR, or STOCH during observing mode. The only hardware injections present in the released data are of the CW type and consist of simulations of 18 spinning neutron stars whose parameters are given on the GWOSC website, [^2] O4a Injections, https://gwosc.org/O4/o4_inj with GW frequencies spanning 12–3000 Hz. Injecting simulated CW signals allows, within the uncertainties of the hardware injection system, a direct end-to-end assessment of degradation in CW signal detection due to imperfect calibration or due to noise subtraction. These injections are almost always present (in O4a, only 1.7% of the released LHO data and 2.6% of LLO data do not have CW injections) but because they are faint and nearly monochromatic they have minimal impact on transient-signal searches.
@@ 4.4 Glitch Subtraction
Table 3: List of O4a events from GWTC:Results with FAR < 1 per year, for which glitch subtraction was performed using BayesWave (Cornish:2020dwh). For each event, we provide the GPS time, the interferometer (IFO) where subtraction was applied, and the time and frequency windows used for subtraction. The time window indicates the aproximate time of the glitch and is referenced from the indicated GPS time, with negative values before the merger and positive values after the merger.
@@ 5.1 Structure of the Default Strain-Data Files
The hdf files contain three folders (or groups): meta, quality and strain.
@@
The gwf files include three channels: one for strain data, one for data quality, and one for injections. The channel names in the GWOSC gwf files are listed in Table 5.
@@ 5.2 Data Quality and Injections in GWOSC Files
The data-quality information is encoded in the default strain release as a bitmask following the structure described in Table 6, with a 1 Hz sampling rate. [^5] See tutorial 3 of https://github.com/gwosc-tutorial/introduction_gwosc_data to learn how to get data-quality information from GWOSC files. The meaning of the data-quality flags is explained in Section 4. For template-based binary coalescence searches and minimally modeled searches, labeled respectively as CBC and BURST in Table 6, the bitmask allows 3 categories of data quality (labeled with CAT1, CAT2 and CAT3). We keep this structure even for runs in which not all the categories have been used. This choice avoids confusion that could arise if the same bit would have different meaning in different runs. A bit value of 1 indicates the data has passed both the indicated level and lower-level checks; for example, a 1 in the CBC_CAT2 bit inidcates the data pass both CBC_CAT1 and CBC_CAT2.
@@ 6 Auxiliary Channels
The O3 and O4a data releases include the set of LIGO auxiliary channels used for noise subtraction and to produce data-quality flags. The auxiliary data release consists of about 40 channels for each LIGO detector for each run. The online documentation includes the name of each available channel along with the associated sampling rate and a short description of the channel content. [^7] See https://gwosc.org/O3/auxiliary/ for O3 and https://gwosc.org/O4/o4a_auxiliary/ for O4a.
@@ 7.1 User Interfaces to the Data
Event Lists are also available to browse specific events using customized queries. Event Lists may include more than one version of an event, as some events have appeared in multiple publications, or have been re-analysed with newer code or models. The Query Page presents the user with a form that allows selection of events based on name, membership in particular releases, various ranges of masses, distance, redshift, effective inspiral spin ($\chi_{\mathrm{eff}}$), $p_{\mathrm{{astro}}}$, SNR, FAR, UTC and GPS times. The query also allows specification of the output format to return, with support for HTML, JSON, CSV and ASCII (see Figure 3).
Figure 3: Event Query Form—A user interface allowing custom selection of events based an selectable attributes.

### v3

@@ 2.1 Observing Time
The instruments operated in a series of observing runs described in GWTC:Introduction, with the following dates:
@@ 4 Data Quality
Noise in the strain data is sometimes approximated as stationary and Gaussian. However, real strain data are also affected by disturbances originating from instrumental and environmental sources (LIGOScientific:2019hgc; LIGO:2024kkz). These artifacts can manifest as broadband, non-Gaussian, short-duration features known as glitches (Nuttall:2018xhi; Glanzer:2022avx), or as narrowband spectral features referred to as spectral lines (Covas_2018). In addition, GW observatories can experience periods of degraded performance due to issues in control systems, calibration procedures, or environmental disturbances. These episodes can compromise data quality over extended durations, reducing the sensitivity of the detectors and potentially biasing astrophysical parameter estimation (Pankow:2018qpo; Macas:2022afm; Kwok:2021zny; Ghonge:2023ksb). To support robust analyses, the LVK employs a suite of data-quality metrics designed to identify and allow mitigation of the impact of instrumental and environmental noise. An overview of LIGO detector characterization efforts is available in LIGO:2024kkz. Detector characterization methods for past observing runs are described in LIGO:2021ppb; Virgo:2022ysc; Nuttall:2015dqa; LIGOScientific:2016gtq; LIGOScientific:2017tza; 2018PhRvD..97h2002C; 2021CQGra..38n5001N; Virgo:2022kwz; and KAGRA:detchar.
@@ 4.1 Hardware Injections
Hardware injection times are available on the GWOSC website via a tool called Timelines [^2] Timelines App, https://gwosc.org/timeline/. Additional notes about hardware injections are also provided in the documentation for each data release. In O4a there were no injections labeled CBC, BURST, DETCHAR, or STOCH during observing mode. The only hardware injections present in the released data are of the CW type and consist of simulations of 18 spinning neutron stars whose parameters are given on the GWOSC website, [^3] O4a Injections, https://gwosc.org/O4/o4_inj with GW frequencies spanning 12–3000 Hz. Injecting simulated CW signals allows, within the uncertainties of the hardware injection system, a direct end-to-end assessment of degradation in CW signal detection due to imperfect calibration or due to noise subtraction. These injections are almost always present (in O4a, only 1.7% of the released LHO data and 2.6% of LLO data do not have CW injections) but because they are faint and nearly monochromatic they have minimal impact on transient-signal searches.
@@ 4.4 Glitch Subtraction
Table 3: List of O4a events from GWTC:Results with FAR < 1 per year, for which glitch subtraction was performed using BayesWave (Cornish:2020dwh). For each event, we provide the GPS time, the interferometer (IFO) where subtraction was applied, and the time and frequency windows used for subtraction. The time window indicates the aproximate time of the glitch and is referenced from the indicated GPS time, with negative values before the merger and positive values after the merger. A similar table is included in GWTC:Results.
@@ 5.1 Structure of the Default Strain-Data Files
The hdf files contain three folders (or groups).
@@
The GWOSC gwf files include three channels (for strain data, data quality, and injections), whose names are listed in Table 5.
@@ 5.2 Data Quality and Injections in GWOSC Files
The data-quality information is encoded in the default strain release as a bitmask following the structure described in Table 6, with a 1 Hz sampling rate. [^6] See tutorial 3 of https://github.com/gwosc-tutorial/introduction_gwosc_data to learn how to get data-quality information from GWOSC files. The meaning of the data-quality flags is explained in Section 4. For template-based binary coalescence searches and minimally modeled searches, labeled respectively as CBC and BURST in Table 6, the bitmask allows 3 categories of data quality (labeled with CAT1, CAT2 and CAT3). We keep this structure even for runs in which not all the categories have been used. This choice avoids confusion that could arise if the same bit would have different meaning in different runs. A bit value of 1 indicates the data has passed both the indicated level and lower-level checks; for example, a 1 in the CBC_CAT2 bit indicates the data pass both CBC_CAT1 and CBC_CAT2.
@@ 6 Auxiliary Channels
The O3 and O4a data releases include the set of LIGO auxiliary channels used for noise subtraction and to produce data-quality flags. These are the ways that auxiliary channels directly contribute to the GWTC-4.0 analysis. The auxiliary data release consists of about 40 channels for each LIGO detector for each run. The online documentation includes the name of each available channel along with the associated sampling rate and a short description of the channel content. [^8] See https://gwosc.org/O3/auxiliary/ for O3 and https://gwosc.org/O4/o4a_auxiliary/ for O4a.
@@ 7.1 User Interfaces to the Data
Event Lists are also available to browse specific events using customized queries. Event Lists may include more than one version of an event, as some events have appeared in multiple publications, or have been re-analysed with newer code or models. The Query Page presents the user with a form that allows selection of events based on name, membership in particular releases, various ranges of masses, distance, redshift, effective inspiral spin ($\chi_{\mathrm{eff}}$), $p_{\mathrm{{astro}}}$, SNR, FAR, UTC and GPS times. The query also allows specification of the output format to return, with support for HTML, JSON, CSV and ASCII.